# Born’s rule from epistemic assumptions

### v1

@@ Abstract
Born’s rule is the recipe for calculating probabilities from quantum mechanical amplitudes. There is no generally accepted derivation of Born’s rule from first principles. In this paper, it is motivated from assumptions that link the ontological content of a proper physical model to the epistemic conditions of the experimental context. More precisely, it is assumed that all knowable distinctions should correspond to distinctions in a proper model. This principle of ontological completeness means, for example, that the probabilistic treatment of the double slit experiment with and without path information should differ. Further, it is assumed that the model should rely only on knowable ontological elements, and that failure to fulfill this principle of ontological minimalism gives rise to wrong predictions. Consequently, probabilities should be assigned only to observable experimental outcomes. Also, the method to calculate such probabilities should not rely on the existence of a precise path of the observed object if this path is not knowable. A similar principle was promoted by Born, even though he did not apply it to probability. Another crucial assumption is that the proper rule to calculate probabilities should be generally valid. It should be applicable in all experimental contexts, regardless the setup that determines which attributes of the studied object are observed, together with the probability to observe each of the associated attribute values. There is no need to refer to the Hilbert space structure of quantum mechanics in the present treatment. Rather, some elements of this structure emerge from the analysis.
@@ 1 Introduction
where $\{A_{l}\}$ is an arbitary set of attributes, and $\{b_{l}\}$ is an arbitrary set of real constants. Assumption (6) was criticized by Grete Hermann [12], and also by John Bell a few decades later [3]. The reason is that for sets of attributes whose values are not simultaneously measurable, like position and momentum, it is not clear what $E(\sum_{l}b_{l}A_{l})$ means, or whether this expression can be defined at all. However, it can be defined in quantum mechanics via the relation $\sum_{l}b_{l}\bar{A}_{l}$ between the Hermitean operators $\{\bar{A}_{l}\}$ that correspond to the attributes $\{A_{l}\}$, giving general meaning to the contested assumption (6). It then then holds in quantum mechanics thanks to the specific form of the rule used to calculate expectation values. However, this rule is closely related to Born’s rule. Therefore the assumptions used by von Neumann are so closely related to the standard postulates for quantum mechanics that the derivation of Born’s rule becomes almost circular.
@@
This starting point in the derivation of Born’s rule is epistemic in nature, in the sense that it stick to relations between probabilities of measurements that are possible to perform. The attitude is similar to that of Werner Heisenberg in his paper from 1925 that pioneered the mathematical formulation of quantum mechanics [11].
@@
In fact, the introduction of agents that make judgments in order to derive Born’s rule means that the nature of both the assumptions and the rule that is derived are operational. The Hilbert space formalism recedes into the abstract background. It may be seen as merely a rug with abstract patterns, on which agents do actual physics, checking probabilities of experimental outcomes.
@@
Figuratively speaking, an attempt is made to pull away the rug of abstract formalism from under the feet of the agents and their experimental arrangement, and to show that no harm is done. Rather, the Hilbert space formalism tentatively reemerges as a generally applicable algebraic representation of the experimental context itself.
The material presented here is a further development of tentative ideas presented in a wider context in previous papers by the same author [17]. This paper is organized as follows. Section 2 introduces concepts and ideas employed in the analysis. The assumptions used to arrive at the desired result are described in section 3. They are divided into epistemic assumptions, presented in section 3.1, and conditions that a generally applicable scheme to calculate probabilities must fulfill, listed in section 3.2. The motivation of Born’s rule is provided in section 4. First, a symbolic formalism that encapsulates the relevant aspects of an experimental context is introduced in section 4.1. Then, it is argued in section 4.2 that a scheme involving Born’s rule is the only acceptable algebraic representation of this formalism, given the assumptions in section 3. Section 5 summarizes the main findings, and section 6 considers the possibility that the chosen epistemic assumptions may be used to gain clarity on more aspects of physical law than Born’s rule.
@@ 2 Ideas and definitions
At the core of the present attempt to motivate Born’s rule is an epistemic approach to physics. It is assumed that the structure of knowledge and the structure of the world reflect each other. A thorough analysis of what can be known about the world and what cannot provides lessons about the world itself. This perspective conditions the choice of definitions and assumptions.
@@ 2.1 Potential knowledge
The physical state of the world is taken to correspond to a state of incomplete knowledge about the world. This state of knowledge will be called the potential knowledge, and will be denoted $PK$. Allowing that potential knowledge may be gained or lost, $PK$ becomes a function of a temporal parameter $n$, that is $PK=PK(n)$. Motivations for the assumption that $PK(n)$ is incomplete at all times $n$ are given in preliminary papers by the present author [17]. Similar hypotheses under the poetic headline ’epistemic horizons’ are reviewed by Szangolies [23].
@@
In the present epistemic approach, an object may simply be defined as an element of perception, which can be subjectively distinguished from other such elements. Knowledge about objects may be defined as proper interpretation of elements of perceptions. A fundamental distinction thus has to be assumed between proper and improper such interpretations. An example would be the distinction between a visual impression that is rooted in objects external to the body, and another such impression that is rooted solely in the brain, like a hallucination or dream. The distinction between knowledge and bare perceptions makes it possible to speak of objects that are hidden at times, such as the sun at night, or microscopic and abstract objects like elementary particles, which are deduced from a systematic analysis of macroscopic observations.
@@
The potential knowledge $PK(n)$ is assumed to be a collective entity that amounts to the union of the potential knowledge at time $n$ of all perceiving subjects. Therefore, a subject must be seen as a fundamental element in the present construction, as well as the existence of several subjects that form a collective. The possibility that different subjects observe the same object is assumed, and may be considered as a definition of the statement that they are part of the same world. The collective nature of $PK(n)$ makes it possible to define a corresponding epistemic physical state that transcends the perspective and perceiving power of individual observers, and thus may become a proper representation of the physical world as a whole.
@@ 2.4 Probabilities and propensities
Karl Popper put forward the propensity interpretation of probability, in which probabilities are seen properties inherent in the physical system that generates the outcome to which these probabilities are assigned [19]. He came to this view thinking about the quantum mechanical double-slit experiment, which convinced him that probabilities must be ’physically real’. To clarify what he meant by this, Popper wrote:
@@ 3.1 Ontological closure
Figure 2: It is assumed that a proper physical model should be complete, meaning that it contains all knowable degrees of freedom. The model should also be minimal, meaning that it does not contain any other degrees of freedom. All degrees of freedom of the model should follow from the forms of appearances, and it should be possible to derive all forms of appearances from the model. Such a model may be called ontologically self-consistent. It is argued that a model that is not ontologically self-consistent is insufficient or wrong.
@@ 3.1.1 Ontological completeness
According to the chosen epistemic ansatz, the knowable world at any time $n$ can be divided into two parts by a moving curtain of ignorance: the part of the world about which we currently have potential knowledge $PK(n)$, and the part of the world about which we have no such knowledge. This distinction should be honored in physical models, meaning that the behaviour of objects or attributes that are part of $PK(n)$ should be modelled differently than those that are not. Specifically, it means that a double-slit experiment $C$ where information about which slit the specimen passes is forever outside potential knowledge should be described and behave differently than a corresponding experimental context $C^{\prime}$ where such path information will become known. In either case, the knowable dynamics corresponds to rules to calculate probabilities $p_{j}$ for different outcomes $A_{j}$. These rules should therefore differ in contexts $C$ and $C^{\prime}$.
The same distinction should be made between the two Mach-Zehnder interferometer contexts in Fig. 1(a) and 1(b). In context (a), with path information, the classical rules of probability apply, giving
@@ 3.1.2 Ontological minimalism
In a paper dedicated to Niels Bohr on the occasion of his 70th birthday [6], Born expands on the subject, using a slightly different wording:
@@ 5.1 Overview
This paper offers a brief review of historical attempts to derive Born’s rule. It is pointed out that it is logically impossible to remove postulates like Born’s rule from quantum mechanics without replacing them with other postulates. However, the new postulates may have a different character than the old ones. A trend is identified where one or several formal postulates, like Born’s rule, are replaced with operational postulates that describe the process of acquiring information by means of experiment, or by epistemic postulates that describe what the experimenter may know about the experiment. The basic idea of this paper is to follow this trend to its natural end, motivating Born’s rule by epistemic assumptions alone, without making any reference to any formal postulates of quantum mechanics.

### v2

@@ Abstract
Born’s rule is the recipe for calculating probabilities from quantum mechanical amplitudes. There is no generally accepted derivation of Born’s rule from first principles. In this paper, it is motivated from assumptions that link the ontological content of a proper physical model to the epistemic conditions of the experimental context. More precisely, it is assumed that all knowable distinctions should correspond to distinctions in a proper model. This principle of ontological completeness means, for example, that the probabilistic treatment of the double slit experiment with and without path information should differ. Further, it is assumed that the model should rely only on knowable ontological elements, and that failure to fulfill this principle of ontological minimalism gives rise to wrong predictions. Consequently, probabilities should be assigned only to observable experimental outcomes. Also, the method to calculate such probabilities should not rely on the existence of a precise path of the observed object if this path is not knowable. A similar principle was promoted by Max Born, even though he did not apply it to probability. Another crucial assumption is that the proper rule to calculate probabilities should be generally valid. It should be applicable in all experimental contexts, regardless the setup that determines which attributes of the studied object are observed, together with the probability to observe each of the associated attribute values. There is no need to refer to the Hilbert space structure of quantum mechanics in the present treatment. Rather, some elements of this structure emerge from the analysis.
@@ 1 Introduction
where $\{A_{l}\}$ is an arbitary set of attributes, and $\{b_{l}\}$ is an arbitrary set of real constants. Assumption (6) was criticized by Grete Hermann [4], and also by John Bell a few decades later [5]. The reason is that for sets of attributes whose values are not simultaneously measurable, like position and momentum, it is not clear what $E(\sum_{l}b_{l}A_{l})$ means, or whether this expression can be defined at all. However, it can be defined in quantum mechanics via the relation $\sum_{l}b_{l}\bar{A}_{l}$ between the Hermitean operators $\{\bar{A}_{l}\}$ that correspond to the attributes $\{A_{l}\}$, giving general meaning to the contested assumption (6). It then holds in quantum mechanics thanks to the specific form of the rule used to calculate expectation values. However, this rule is closely related to Born’s rule. Therefore the assumptions used by von Neumann are so closely related to the standard postulates for quantum mechanics that the derivation of Born’s rule becomes virtually circular.
@@
This starting point in the derivation of Born’s rule is epistemic in nature, in the sense that it sticks to relations between probabilities of measurements that are possible to perform. The attitude is similar to that of Werner Heisenberg in his paper from 1925 that pioneered the mathematical formulation of quantum mechanics [10].
@@
In fact, the introduction of agents that make judgments in order to derive Born’s rule means that the nature of both the assumptions and the rule that is derived are operational. The Hilbert space formalism recedes into the abstract background. It may be seen as merely a carpet with abstract patterns, on which agents do actual physics, checking probabilities of experimental outcomes.
@@
Figuratively speaking, an attempt is made to pull away the carpet of abstract formalism from under the feet of the agents and their experimental arrangement, and to show that no harm is done. Rather, the Hilbert space formalism tentatively reemerges as a generally applicable algebraic representation of the experimental context itself.
The material presented here is a further development of tentative ideas presented in a wider context in previous papers by the same author [15]. The present paper is organized as follows. Section 2 introduces concepts and ideas employed in the analysis. The assumptions used to arrive at the desired result are described in section 3. They are divided into epistemic assumptions, presented in section 3.1, and conditions that a generally applicable scheme to calculate probabilities must fulfill, listed in section 3.2. The motivation of Born’s rule is provided in section 4. First, a symbolic formalism that encapsulates the relevant aspects of an experimental context is introduced in section 4.1. Then, it is argued in section 4.2 that a scheme involving Born’s rule is the only acceptable algebraic representation of this formalism, given the assumptions in section 3. Section 5 summarizes the main findings, and section 6 considers the possibility that the chosen epistemic assumptions may be used to gain clarity on more aspects of physical law than Born’s rule.
@@ 2 Ideas and definitions
At the core of the present attempt to motivate Born’s rule is an epistemic approach to physics. It is assumed that the structure of knowledge and the structure of the physical world reflect each other. A thorough analysis of what can be known about this world and what cannot provides lessons about the world itself. This perspective conditions the choice of definitions and assumptions.
Perceiving and knowing subjects are necessary ingredients in such a world view. The relation between the subjective and the objective aspects of the world becomes fundamental, as illustrated in Fig. 1. It is considered meaningless to talk about perception in itself without any objects to perceive, just as it is empirically meaningless to consider objects that cannot potentially be perceived. The subjective and objective aspects of the world are thus seen as two sides of a coin: distinguishable but inseparable. This model is similar to John Wheeler’s idea of a participatory universe [16].
To be able to speak about and identify objective physical law, like Born’s rule, we have to be able to interpret and generalize our subjective perceptions. A fundamental distinction between proper and improper such interpretations has to be assumed. This distinction transcends the perceptions themselves. In this picture, knowledge corresponds to properly interpreted perceptions. To exemplify, a dream is properly interpreted as such, whereas an improper interpretation would be that the perceptions during the dream correspond to objects external to the brain of the dreamer.
Figure 1: Schematic illustration of the world view that inspires the assumptions used to motivate Born’s rule. In this view, there is nothing objective in the physical world independent of the subjective, and the subjective is completely embedded in the objective in the sense that all perceptions can be correlated with physical states of bodies of perceiving subjects.
@@ 2.1 Potential knowledge
The physical state of the world is taken to correspond to a state of knowledge called the potential knowledge, which will be denoted $PK$. The aim of the present discussion about potential knowledge $PK$ is to provide the necessary conceptual background to the use of $PK$ in the following analysis. A more detailed discussion is found in Ref. [15].
Allowing for potential knowledge to be gained or lost, $PK$ becomes a function of a temporal parameter $n$, that is $PK=PK(n)$. Potential knowledge $PK(n)$ is assumed to be incomplete at all times $n$. The reason is intuitively clear from Figure 1. This knowledge is represented in the physical state of the bodies of all perceiving subjects, and the set of these bodies is considered to be a proper subset of the physical world. Thus, if $PK(n)$ were complete, this knowledge would be represented in a proper subset of itself. This is not possible under the reasonable assumption that the states of the bodies and the state of the external world can vary independently. Similar hypotheses about the incompleteness of knowledge are reviewed by Szangolies under the poetic headline epistemic horizons [17].
@@
In the present epistemic approach, an object may simply be defined as an element of perception, which can be subjectively distinguished from other such elements. Knowledge about objects may be defined as proper interpretation of elements of perceptions. The distinction between knowledge and sense perceptions makes it possible to speak of objects that are hidden at times, such as the sun at night, or microscopic and abstract objects like elementary particles, which are deduced from a systematic analysis of macroscopic observations.
@@
The existence of several subjects is assumed, and the potential knowledge $PK(n)$ is assumed to be a collective entity that amounts to the union of the potential knowledge at time $n$ of all perceiving subjects. The possibility is assumed that different subjects observe the same object, that is, that they share elements of perception. This circumstance may be considered as a definition of the statement that these subjects are part of the same world. The collective nature of $PK(n)$ makes it possible to define a corresponding epistemic physical state that transcends the perspective and perceiving power of individual observers, and thus may become a proper representation of the physical world as a whole.
The underlying assumption that different subjects may share perceptions is seen as fundamental, meaning that it cannot be fully accounted for in terms of something else. The most common attempted such explanation is that hypothesized observer-independent objects emit electromagnetic radiation that enters the eyes of different observers triggering a common perception. Even though such observer-independent objects are at odds with the world view proposed here, the present model provides objectivity to perceptions in the sense that they transcend the individual perceiving subject when they are shared, just as the proper interpretation of such shared perceptions is assumed to transcend the individual interpreting subject, allowing for common knowledge.
It should be stressed that potential knowledge at the conceptual level is considered to be independent of any particular representation or encoding of this knowledge, such as formal propositions or strings of bits. The possibility to distinguish knowledge from its representations follows from the assumption that subjective perceptions and insights about these perceptions are fundamental notions.
The epistemic physical state $S(n)$ introduced in Ref. [15] is an example of a representation of $PK(n)$. This state is not needed in the present study, but it can be seen as a generalisation of the set $S_{A}$ to be discussed in the next section.
@@ 2.4 Probabilities and propensities
Karl Popper put forward the propensity interpretation of probability, in which probabilities are seen as properties inherent in the physical system that generates the outcome to which these probabilities are assigned [22]. He came to this view thinking about the quantum mechanical double-slit experiment, which convinced him that probabilities must be ’physically real’. To clarify what he meant by this, Popper wrote:
@@ 3 Assumptions
The philosophical ideas introduced in section 2 are presented in as much detail as necessary to make it understandable why the author of this paper has chosen the following assumptions, but not in any more detail than that. These assumptions provide the foundation of the present study, not the underlying philosophy that has inspired them.
@@ 3.1 Ontological closure
Figure 3: It is assumed that a proper physical model should be complete, meaning that it contains all knowable degrees of freedom. The model should also be minimal, meaning that it does not contain any other degrees of freedom. All degrees of freedom of the model should follow from the forms of appearances, and it should be possible to derive all forms of appearances from the model. Such a model may be called ontologically self-consistent. It is argued that a model that is not ontologically self-consistent is insufficient or wrong. Such a loop of self-consisteny reflects the loop that describes the relation between the subjective and objective aspects of the world according to Fig. 1.
@@ 3.1.1 Ontological completeness
According to the chosen epistemic ansatz, the knowable world at any time $n$ can be divided into two parts by a moving curtain of ignorance: the part of the world about which we currently have potential knowledge $PK(n)$, and the part of the world about which we have no such knowledge. This distinction should be honored in physical models, meaning that the behaviour of objects or attributes that are part of $PK(n)$ should be modelled differently than those that are not. Specifically, it means that a double-slit experiment $C$ where information about which slit the specimen passes is forever outside potential knowledge should be described and behave differently than a corresponding experimental context $C^{\prime}$ where such path information will become known. In either case, the knowable dynamics corresponds to procedures to calculate probabilities $p_{j}$ for different outcomes $A_{j}$. These procedures should therefore differ in contexts $C$ and $C^{\prime}$.
The same distinction should be made between the two Mach-Zehnder interferometer contexts in Fig. 2(a) and 2(b). In context (a), with path information, the classical rules of probability apply, giving
@@ 3.1.2 Ontological minimalism
In a paper dedicated to Niels Bohr on the occasion of his 70th birthday [27], Born expands on the subject:
@@ 5.1 Overview
This paper offers a brief review of previous attempts to derive Born’s rule. It is pointed out that it is logically impossible to remove postulates like Born’s rule from quantum mechanics without replacing them with other postulates. However, the new postulates may have a different character than the old ones. A trend is identified where one or several formal postulates, like Born’s rule, are replaced with operational postulates that describe the process of acquiring information by means of experiment, or by epistemic postulates that describe what the experimenter may know about the experiment. The basic idea of this paper is to follow this trend to its natural end, motivating Born’s rule by epistemic assumptions alone, without making any reference to any formal postulates of quantum mechanics.